OpenStudy (anonymous):

I will give a medal if you help or fan. @ganeshie8 @undeadknight26 @xapproachesinfinity @SolomonZelman

OpenStudy (anonymous):

Which of the following statements best describes the graph of 3x – 2y = 4? It is a straight line joining the points (6, 2), (5, 1), and (7, 3). It is a straight line joining the points (0, –2), (2, 1), and (–2, –5). It is a curve joining the points (3, –2), (2, 3), and (4, 1). It is a curve joining the points (0, 2), (–2, –1), and (4, 1).

OpenStudy (anonymous):

can you put this in y=mx+b form

OpenStudy (solomonzelman):

Well, it is definitely not a curve :)

undeadknight26 (undeadknight26):

Ill give ya its a straight line.

OpenStudy (solomonzelman):

I have enough medals and fans... man, just plug in your points from option A, and then points from option B .

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

B

OpenStudy (solomonzelman):

Hint: y-intercept

OpenStudy (solomonzelman):

YEs

OpenStudy (solomonzelman):

It is B

OpenStudy (anonymous):

Do you mind helping with more

OpenStudy (solomonzelman):

I can try at least.

OpenStudy (anonymous):

Just to check my work.

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

The tables below show the values of f(x) and g(x) for different values of x: f(x) = 2(3)x x f(x) -2 0.22 -1 0.67 0 2 1 6 2 18 g(x) = 3x + 3 x g(x) -2 3.11 -1 3.33 0 4 1 6 2 12 Based on the tables, what is the solution to the equation 2(3)x = 3x + 3? x = –1 x = 0 x = 1 x = 6

OpenStudy (solomonzelman):

I don;t really get what `2(3)x` is supposed to mean

OpenStudy (solomonzelman):

2(3)^x ?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I think its C

OpenStudy (solomonzelman):

The y-coordinate is approximately 1.

OpenStudy (solomonzelman):

the x-coordinate is about -1, well -0.686

OpenStudy (solomonzelman):

I graphed the functions that is where they intersect.

OpenStudy (anonymous):

so i was wrong?

OpenStudy (anonymous):

It is x=1. But next question.

OpenStudy (anonymous):

The two lines, X and Y, are graphed below: Line X is drawn by joining ordered pairs negative3,12 and 7,negative16. Line Y is drawn by joining ordered pairs 0, negative14 and 11, 8 Determine the solution and the reasoning that justifies the solution to the systems of equations. (2, 7), because this point is true for both the equations (4, –6), because this point lies only on one of the two lines (4, –6), because this point makes both the equations true (2, 7), because the lines intersect the x-axis at these points @SolomonZelman

OpenStudy (anonymous):

@phi can you look over this?

OpenStudy (anonymous):

B for the 3rd^

OpenStudy (phi):

which question?

OpenStudy (anonymous):

the one i just posted.

OpenStudy (phi):

based on the wording B and D are not the answer. note for B: (4, –6), because this point lies only on *one* of the two lines the solution must lie on *both lines*. so B is incorrect so it is either A or C

OpenStudy (phi):

It is hard to read numbers mixed with words. are the points (-3,12) and (7,-16) for the first line? and (0,-14) , (11,8) for the points on the second line ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

There seems to be something wrong. we can find the equation of the second line to be y= 2x -14 and of the two choices (2,7) and (4,-6) on (4,-6) is on this line in other words, -6 = 2*4-14 is true however, the equation of the first line does not contain either of the two choices.

OpenStudy (anonymous):

i will skip this one and come back to it.

OpenStudy (phi):

I think they have a typo in the points given for the first line

OpenStudy (anonymous):

An equation is shown below: 3 over 2 multiplied by x plus 7 over 2 equals 2 to the power of x

OpenStudy (anonymous):

i got 1

OpenStudy (phi):

Here is the graph. Notice that (6,-4) is on the 2nd line and close to the first. point (7,2) (the red dot) is not on either line.

OpenStudy (anonymous):

no 3

OpenStudy (phi):

can you use the equation editor?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

what is that?

OpenStudy (phi):

It's the button on the bottom left of the input area it lets you put in complicated equations

OpenStudy (phi):

but this is my guess \[ \frac{3x}{2} + \frac{7}{2} = 2^x \] if you multiply both sides by 2 you get \[ 3x + 7 = 2^{x+1} \] this is a difficult equation to solve. But you can test each choice, and see if the works.

OpenStudy (phi):

x=3 gives 9+7 = 2^(3+1) 16= 2^4 16=16 which is true. so x=3 is an answer.

OpenStudy (anonymous):

yes i did it on paper and got x=3 but i subtracted the first time and got 1.

OpenStudy (anonymous):

The graph below shows two functions: function f of x is a straight line which joins the ordered pairs negative 2.8,16 and 3.5, negative 2.5. Function g of x is a curved line which joins the ordered pairs negative 1.9, 13.5 and 2.6, 0 Based on the graph, what are the approximate solutions to the equation –3x + 8 = (0.25)x? 1.9 and –2.6 8 and 2.6 1 and 2.6 –1.9 and 2.6

OpenStudy (anonymous):

I have no idea about this one ^

OpenStudy (phi):

you look for where the two curves cross each other your graph should look like this

OpenStudy (phi):

they are asking for the x values of the two points where the lines cross. do you see the lines cross in 2 places?

OpenStudy (anonymous):

Its on 2.7 though

OpenStudy (phi):

yes, but they say approximate so we have some wiggle room on this one.

OpenStudy (anonymous):

D

OpenStudy (phi):

yes. if we try for example x = -1.9 in the equation it should (almost) work –3x + 8 = (0.25)^x 13.7 =? 13.9 (according to my calculator) which is pretty close (we would need a lot more decimals to make it an exact match)

OpenStudy (anonymous):

so i round up?

OpenStudy (anonymous):

for a whole number

OpenStudy (phi):

no, I was just checking one of their answers to show it is reasonable. when we read off answers from a graph it is only going to be "close" unless we zoom in. so D is the answer